OpenStudy (anonymous):

the foci of ((x-5)^2 /25) + ((y+3)^2 /9) = 1

OpenStudy (anonymous):

(1, -3) | (9, -3)

OpenStudy (anonymous):

Does that answer your question?

OpenStudy (apoorvk):

Okay can you guess what changes "x-5" and "y+3" cause in \[\frac{(x-5)^2}{25} + \frac{(y+3)^2}{9} = 1\] I mean what is their significance? Because the equation of a standard ellipse is : \[\frac{x^2}{a^2} + \frac{y^2}{b^2} = 1\] Right? So what changes are those incorporating into this? Try to think on the lines of a circle's equation if you've studied them...

OpenStudy (anonymous):

thanks

OpenStudy (apoorvk):

Hmm, already understood that?

OpenStudy (anonymous):

yes..i think i do ..

OpenStudy (apoorvk):

Brilliant then! still for my own satisfaction, can you answer my question above?